OpenStudy (anonymous):

A=1/2h(b1+b2) solve for b2 Help please!

OpenStudy (anonymous):

A=1/2h(b1+b2) 2A/h=(b1+b2) b2=2A/h-b1

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

YW